# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ Abstract
We propose both a theoretical and a methodological framework to address a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we present a solution by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove, establishing a strong foundation for regressing complex SO(3)-equivariant targets. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm. Our method boosts Hamiltonian prediction accuracy by up to 40% and enhances downstream physical quantities, such as occupied orbital energy, by a maximum of 76%. Our method also significantly promotes the acceleration ratios for the convergence of traditional Density Functional Theory (DFT) methods.
@@ 1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research (Zhang et al., 2023). For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ (Unke et al., 2021; Gu et al., 2022; Li et al., 2022; Yu et al., 2023b; Gong et al., 2023), opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design, as well as groundbreaking molecular pharmaceutical research that was previously unimaginable.
@@
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, it is challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on modeling physical systems. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. This issue is commonly found in physics-oriented machine learning tasks that demand both strict equivariance and fine-grained generalization performance, as analyzed by Zitnick et al. (2022). Several recent efforts have been made to alleviate this issue (Zitnick et al., 2022; Passaro & Zitnick, 2023; Wang et al., 2024c; Yin et al., 2024), however, they often compromised strict SO(3)-equivariance.
@@
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity (Resnick, 1991), the spacetime interval between events is invariant under Lorentz transformations, and is fundamental in formulating physical laws that involve equivariant quantities. For another example, the total electronic energy of an atomic system is SO(3)-invariant, while the derived force fields are SO(3)-equivariant. We aim to extend the relationship between invariance and equivariance from specific physical quantities to a more general representation learning framework. From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones. Specifically:
@@
First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{Q}$ denotes the SO(3)-equivariant regression targets we aim to predict. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ in the training phase. Taking SO(3)-invariant features as a bridge, this mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove, laying a solid foundation for accurately inferring complex SO(3)-equivariant targets.
@@
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deep understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as pointed out by Yin et al. (2024). Our method significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a). It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories. Furthermore, as observed from the experiments on the QH9 benchmark series, our approach also substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonian including occupied orbital energy and electronic wavefunction. Moreover, our method also demonstrates superior performance in accelerating the convergence of classical DFT by providing predicted Hamiltonians as initialization matrices. Our leading performance comprehensively demonstrates that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
@@ 2 Related Work
The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based symmetry operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features. These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), molecular property prediction and dynamic simulation (Musaelian et al., 2023; Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023). However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l$ greater than one. This restriction severely limits the network’s capability to model complex non-linear mappings. To alleviate this issue, methods like DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023b) and Equiformer (Liao & Smidt, 2023) introduced a gated activation mechanism that feeds SO(3)-invariant features ($l=0$) into non-linear activation functions and uses these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. Nevertheless, multiplying SO(3)-equivariant features with linear coefficients may not fundamentally improve their non-linear expressiveness.
@@
In order to improve non-linear expressiveness, Zitnick et al. (2022) decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions. These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions. In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b; c), this approach has demonstrated a remarkable capacity to fit complex functions. However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al. (2022) proposed a local coordinate strategy, projecting rotating global coordinates onto SO(3)-invariant local ones for the non-linear neural network to encode. However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al. (2024) proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians. In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results. In this paper, we propose a framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks to resolve the equivariance-expressiveness dilemma.
3 Problem Formalization
For a background introduction on groups, group representations, equivariance, and invariance, please refer to Appendix A. Here, we directly focus on the equivariance of physical quantities under 3D rotational operations that form the SO(3) group. Let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}={\bf{D}}^{l_{p}}({\bf{R}})\cdot{\bf{Q}}^{l_{p}\otimes l_{q}}\cdot{({\bf{D}}^{l_{q}}({\bf{R}}))^{\dagger}}$ (1)
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively; ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. 1 but with a simpler form:
${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form. To obey the equivariant law of Eq. 2 for the regression target, these hidden representations must also satisfy the same form of equivariance:
${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonian in this context, is the core problem this paper aims at solving.

### v5

# TraceGrad: a Framework Learning Expressive SO(3)-equivariant Non-linear Representations for Electronic-Structure Hamiltonian Prediction

@@ Abstract
We propose a framework to combine strong non-linear expressiveness with strict SO(3)-equivariance in prediction of the electronic-structure Hamiltonian, by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. The proposed framework, called TraceGrad, first constructs theoretical SO(3)-invariant trace quantities derived from the Hamiltonian targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the learning of invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate powerful non-linear expressive capabilities into SO(3)-equivariant features with consistency of physical dimensions to the regression targets, while theoretically preserving equivariant properties, establishing a strong foundation for predicting Hamiltonian. Our method achieves state-of-the-art performance in prediction accuracy across eight challenging benchmark databases on Hamiltonian prediction. Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration performance for the traditional Density Functional Theory algorithms.
Electronic-structure Hamiltonian prediction; SO(3)-equivariant representation learning; Non-linear expressiveness
@@ 1 Introduction
Electronic structure calculations provide crucial insights into the behavior of electrons in condensed matter, forming the foundation for predicting material properties such as conductivity, magnetism, optical response, and chemical reactivity, playing a key role in applications ranging from semiconductors to renewable energy and catalysis, ultimately helping to reshape our lives. The core of electronic structure calculations lies in solving the Hamiltonian matrices, from which critical physical quantities such as orbital energy, band structure and electronic wavefunction are induced. However, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from the extremely time-consuming self-consistent field (SCF) algorithms used to obtain the Hamiltonians. These algorithms involve the exhaustive iterative diagonalization of large matrices, with each diagonalization scaling with a high computational complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system. With advantages in computational complexity and generalization capabilities, the deep learning paradigm has significantly advanced physics research (Zhang et al., 2023). In the task of predicting the Hamiltonians, deep learning approaches (Unke et al., 2021; Yu et al., 2023b; Gong et al., 2023; Zhang et al., 2024; Wang et al., 2024c) bypass the SCF process, dramatically reducing the computational complexity of solving Hamiltonians, opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design that was previously unimaginable. See Appendix B for more details.
@@
To align with fundamental physical laws, deep learning methods for Hamiltonian prediction must strictly adhere to equivariance under 3D rotational transformations, i.e., elements from the SO(3) group. However, it remains challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on predicting Hamiltonians. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. This issue is commonly encountered in physics-oriented machine learning tasks that require both strict equivariance and fine-grained generalization performance, as discussed by Zitnick et al. (2022). Several recent methods, such as the gated mechanism (Weiler et al., 2018), spherical channel-based approaches (Zitnick et al., 2022; Liao et al., 2024; Wang et al., 2024c), local coordinate strategies (Li et al., 2022), and hybrid regression methods (Yin et al., 2024), have been proposed to address this challenge. However, combining strong non-linear expressiveness with strict SO(3)-equivariance remains challenging for the Hamiltonian prediction task, as analyzed in detail in Section 2.
@@
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning for electronic-structure Hamiltonian prediction. We are inspired by the insight that invariant quantities in transformation (Resnick, 1991) often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties, and aim to extend the relationship between invariance and equivariance from specific physical quantities to hidden representations of neural networks. From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to supervising high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones. Specifically:
@@
First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{H}$ denotes the SO(3)-equivariant regression targets, i.e., the block of Hamiltonian matrix. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{H}$ without requiring additional labeling resources.
@@
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{H}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ in the training phase. Taking SO(3)-invariant features as a bridge, this mechanism can incorporate powerful non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove. Compared to the gated mechanism (Weiler et al., 2018), the proposed mechanism naturally follows the correspondence of physical dimensions between the SO(3)-equivariant Hamiltonian and the SO(3)-invariant trace quantity, laying a solid foundation for regressing Hamiltonian with the help of supervision signals from the trace quantity.
@@
Experimental results demonstrate that, our TraceGrad framework significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a). It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories. Furthermore, as observed from the experiments on the QH9 benchmark series, our approach also substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonian including occupied orbital energy and electronic wavefunction. Moreover, our method also demonstrates superior performance in accelerating the convergence of classical DFT by providing predicted Hamiltonians as initialization matrices. Our leading performance comprehensively demonstrates that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
@@ 2 Related Work
The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based tensor operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features. These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), energy and force prediction (Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023). However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l\geq 1$. This restriction severely limits the network’s capability to model complex non-linear mappings. To alleviate this issue, Hamiltonian prediction methods like DeepH-E3 (Gong et al., 2023) and QHNet (Yu et al., 2023b) introduced a gated activation mechanism (Weiler et al., 2018) that fed SO(3)-invariant features ($l=0$) into non-linear activation functions and used these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. However, these methods only indirectly supervised the learning of SO(3)-invariant features through SO(3)-equivariant Hamiltonians taking SO(3)-equivariant representations as a bridge, and the lack of direct SO(3)-invariant supervision signals may result in insufficient learning of SO(3)-invariant features, limiting their expressiveness to assist SO(3)-equivariant features. Furthermore, the gated mechanism, which directly multiplies SO(3)-equivariant features with SO(3)-invariant features to yield new SO(3)-equivariant features, lacks consideration of the physical dimensional relationship between the actual SO(3)-equivariant and SO(3)-invariant physical quantities, making it difficult to achieve a perfect coupling of their representations. See Remark 4.3 for a more comprehensive analysis.
@@
In order to improve non-linear expressiveness, Zitnick et al. (2022) proposed spherical channel methods, which decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions. These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions. In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b, c), this approach has demonstrated a remarkable capacity to fit complex functions. However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al. (2022) proposed a local coordinate strategy, projecting rotating global coordinates onto SO(3)-invariant local ones for the non-linear neural network to encode. However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al. (2024) proposed a hybrid regression framework, where the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation. In many applications of electronic structure calculations, the demands for symmetry are extremely high. Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results. In this paper, we seek for combining strict SO(3)-equivariance with the powerful non-linear neural expressiveness to resolve the equivariance-expressiveness dilemma for electronic-structure Hamiltonian prediction.
3 Preliminary
In order to better understand our theory and methods, we recommend that the readers first refer to the preliminary in Appendices: Appendix A provides the definitions of foundational concepts relevant to the research topic; Appendix B describes the application tasks, specifically the electronic-structure Hamiltonian prediction task; Appendix C formalizes Hamiltonian prediction as an SO(3)-equivariant non-linear representation learning problem.